OpenStudy (anonymous):

Solve the following system. (Use (x, y) format) x2 + y2 = 25 y2 - x2 = 7

OpenStudy (anonymous):

There are several ways to solve this. Let's try substituting. We know that in a system of equations, all the equations use the same variables. That means that the x2 from one system is the same as the x2 from the other system--that's an important detail though. Is it supposed to be x2 as in x*2 or x2 as in x with the subscript of 2? Usually in algebra, we write the coefficient(the number on the variable) first. Anyway, solve for y in terms of x in one of the equations and then substitute it back into the other equation. I will help you with this if you have additional trouble.

OpenStudy (anonymous):

I understand. Thank you!

OpenStudy (anonymous):

One of the most common mistakes when solving these types of problems is forgetting to distribute. For example: x+y=25 -x+y=7 x=25-y this is the key step. You should always substitute with parentheses, to make it clear to your teacher what you're doing, and to make it easier to avoid mistakes. -(25-y)=7 -25+y=7 y=7+25=32